OpenStudy (anonymous):

is the square route of 5/8 rational or irrational?

OpenStudy (anonymous):

Square roots are only rational if they are being taken of perfect squares. Do you know how to determine if something is a perfect square?

OpenStudy (anonymous):

\[\sqrt{\frac{5}{8}}\] \[\frac{ \sqrt{5} }{ 2\sqrt{2} }\] both the square root terms are irrational. so their quotient will also be irrational.